OpenStudy (alyssajobug):

If a sound wave has a specific intensity of 6 μW/m^2 and propagates symmetrically over a spherical radius of 4 m, what is the power of the sound wave?

OpenStudy (bmk614):

\[I=\frac{ P }{ 4 \pi r^2 }\] Plug in 4 for r and 6 for I. Solve for P.

OpenStudy (alyssajobug):

So the answer would be 1,205.76?

OpenStudy (alyssajobug):

@bmk614

OpenStudy (alyssajobug):

6=P/4(3.14)(16) 6=P/200.96 6 x 200.96=P 1,205.76=P

OpenStudy (bmk614):

Yes. That is what I got!

OpenStudy (alyssajobug):

Perfect, thank you so much for all of your help!

OpenStudy (alyssajobug):

What would the label be for that though? 1,205.76 μW?

OpenStudy (bmk614):

Yes! That is correct!